OpenStudy (anonymous):

Is the below line from this equation parallel and why or why not #38: x1 y1 x2 y2 line1 k passes through (4, 2) and (-4, 4) x1 y1 x2 y2 line2 h passes through (-8, 1) and (8, -3) Is line k < h? mk = 4 – 2 = 2 = 1 -4 -4 -8 -4 mh = -3 – 1 = -4 = -1 8 – (-8) 16 4

OpenStudy (anonymous):

To answer this question, let's find the slopes given by: \[m = \frac{ (y2-y1) }{ (x2-x1) }\] line k => mk = (4-2)/(-4-4) = 2/-8 = -1/4 line h => mh = (-3-1)/(8--8)=-4/16 = -1/4 m1=m2 Since these two lines have identical slopes, then these lines are parallel.

OpenStudy (anonymous):

Thank you so much. I can't tell you how much I appreciate your answering my question. Now, another quick question ... someone emailed me back and threatened me not to use other forms of communication on this site if math is the only subject I need help in. Sorry, never used this site before. I guess I went into the wrong tab or something. Will be needing help here quite a bit. Exactly what steps do I take because I don't even know how I got to you. Was just putting the question out there. Please advise. Thanks.

OpenStudy (anonymous):

@stamps3 you are welcome:) Actually I don't know who has threatened u via email. If the message is from notifier@openstudy.com, then probably u have violate the openstudy's code of conduct [ Read it here http://openstudy.com/code-of-conduct ] If the message is from some other person I don't think u have to worry about that. This site is promoting learning, so feel free to ask question from any of the subjects u are studying. I'll be here to help u:) ~Happy Studying!:)